Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Fiori, Tringali - 17:59 Saturday 16 December 2023 (62759) Print this report
ENV shift - DET and cryo-traps tappings and acoustic injections CEB

The shift was dedicated to further investigation of scattered light at the DET viewports and minilinks. We also performed tappings at the DET and INJ cryo traps. Finally we performed acoustic injections in the CEB in order to improve the measured coupling function.

DETECTION lab

We installed one accelerometer on SDB2 flange, oriented along the beam (Figure 2). We perfomed excitations of all the viewports and minilink pipes by means of the BlueTooth shaker. The scheme (Figure 1) shows the location of the viewports. Figures from 2 to 10 show the position of the shaker during each injection. We injected colored noise between 20 Hz and 1 kHz (but effectively the shaker did not reproduce vibrations below 40Hz) and with the same volume everywhere.

The details of the actions are reported in the attached log file.

Some results:

  • B1 viewport - peaks around 66 Hz, associated to resonance of SDB2 chamber or the rectangular flange (Figure 11)
  • HWS viewport - similar excitation of SDB2 (accelerometer) peaking at 60 Hz and 66Hz (Figure 12)
  • B1p viewport - we are exciting mainly SDB1, several peaks appear in hrec, 61 Hz, 66 Hz (now smaller, 184,281, 304, 321, 500, 504, 578, 595. 614 (Figure 13, Figure 14, Figure 15)
  • B5 viewport - hrec peaks beween 480 and 650 Hz are excited the most (some excitation again at 66 Hz and 63 Hz: they seem to be associated resp. to mechanical modes of SDB2 (66) and SDB1 (63)) (Figure 16)
  • B5 pipe - we notice large difference respet to B5 viewport position: now almost NO PEAK EXCITED (apart from 66Hz, again associated to large excitation of SDB2 side at this frequency) and 860 Hz. (Figure 17)
  • HWS pipe - no effect in hrec (almost no seism transmitted to SDB1) only 66Hz (associated to 66Hz seism on SDB2) (Figure 18)
  • B1p pipe - effect similar to "B1p viewport" location, this is reasonable since the shaker is close to the viewport with no bellow in between. (Figure 19)
  • B1p  pipe - We repeated with the shaker on the pipe section closer to SDB2. This time we excited mostly SDB2 and in hrec only peaks around 60 Hz: percisely: 53 Hz, 59Hz, 66 Hz (Figure 20)
  • B1 pipe - nothing in hrec (apart small 66 Hz) (Figure 21)
  • We injected a set of lines from 307 Hz to 312 Hz to check if we could excite the 309.5 Hz bump in Hrec. We see that 309-310 Hz region ("Maddalena's" bump in hrec) responds with amplified excitation in hrec and two bumps at +- 294.7 and 320.5 Hz (baricenter at 307.6 Hz) (Figure 22)

DET cryo trap

  • We tapped 5 locations (two sides, two locations on the central part - belly - of the CT, and at the bellow). We excited a little bit peaks at 500 Hz and 525 Hz.

INJ cryo trap

  • We tapped in similar locations. No apparent effect in hrec. Data to be better examined.

Acoustic injections CEB

  • We injected colored noise between 10 and 1190 Hz with the loudspeaker in two locations: DET terrace (Figure 23) and small terrace at the back of the DAQ room.
  • We injected sweep line between 10 Hz and 2000 Hz (300s periodicity) and 0.04 V in DET terrace.
Images attached to this report
Non-image files attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×